OpenStudy (anonymous):

Which bond to invest in: Bond A has a Face Value of $500,000, life of 4 years, coupon rate of 6% p.a. And a yield return of 5% p.a. Bond B has a Face Value of $100,000, life of 2 years, coupon rate of 7% p.a. And a yield return of 6% p.a

OpenStudy (anonymous):

Could you please help me @satellite73

OpenStudy (anonymous):

Am I supposed to use the formula PV = FV (1+i)^-n?

OpenStudy (anonymous):

I need help please @agent0smith

OpenStudy (anonymous):

@Nurali

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@ash2326

OpenStudy (tkhunny):

Bond A has a Face Value of $500,000, life of 4 years, coupon rate of 6% p.a. And a yield return of 5% p.a. Coupon = 500000 * 0.06/2 = 500000 * 0.03 = 15000 (semi-annual coupon) Yield = 0.05 Semi-Annual Yield = 0.05/2 = 0.025 v = 1/(1+0.025) 15000(v + v^2 + ... + v^8) + 500000v^8 Can you add all that up?

OpenStudy (tkhunny):

Generally, I'm a little worried about this question. This is a very poor choice. Bond B lasts only 2 years. What will you be doing with the cash after 2 years? It appears that this is not well-defined. The choice, as presented, is really this: Bond A for 4 years. Bond B for 2 years, then stuff the money in a mattress for 2 years. It's pretty easy to make a choice, here. No calculation required.

OpenStudy (anonymous):

How do you add them all up? I might have a different way to do it. Give me a sec

OpenStudy (anonymous):

In addition why did you convert to semi annual?

OpenStudy (anonymous):

@tkhunny

OpenStudy (anonymous):

@mathstudent55

OpenStudy (anonymous):

I guess nobody knows how to solve it including myself :)

OpenStudy (anonymous):

I need your help @hero. This question has been hanging around for 8 hours. I have an exam on Friday. Could you please help

OpenStudy (tkhunny):

Coupons are normally assumed to be semi-annual. A form with which you should be familiar is the sum of a geometric series, finite or infinite. 15000(v + v^2 + ... + v^8) + 500000v^8 = \(15000\dfrac{v-v^{9}}{1-v} + 500000v^{8}\) You do B.